OpenStudy (anonymous):

After how many seconds would Black Dragon and Wild Thing have the same amount of fuel left in their tanks? Round your answer to the nearest hundredth of a second. (Enter only the number in the form 0.123) Hint: Select the two relevant equations from the last problem and solve them for the variable s. seconds

OpenStudy (anonymous):

Black Dragon and Red Rocket each have 18-gallon fuel tanks. Wild Thing and Lucky Lady have 20-gallon fuel tanks. Match the linear equation that could be used to model the fuel remaining, f, in each car after t seconds. 1. Wild Thing A. f(s) = 18 - 2.14s 2. Red Rocket B. f(s) = 20 - 2.68s 3. Lucky Lady C. f(s) = 20 - 3.06s 4. Black Dragon D. f(s) = 18 - 2.33s Those are the correct answers for the graph and it is needed for my original question i am lost on this problem please help me thankyou

OpenStudy (zpupster):

if you divide fuel used by time and match the (s) you will get your answer. ie black dragon: 14.2/6.1 =2.33 you see 4 is D and so on and so on....

OpenStudy (anonymous):

@zpupster i already solved the 1-4 i need the answer for the first question it requires the graph to be answered.

OpenStudy (missmob):

black dragon

OpenStudy (anonymous):

@missmob wat is that suppose to mean?? that doesnt answer anything

OpenStudy (zpupster):

following the formula I gave S: 2 is A 1 is c 3 is b and 4 is d........... Match the linear equation